rate and unit rate
A 50-lb bag of Glossy Coat Horse Feed costs $23.50. A 25-lb bag costs $15.50. How much money per pound would you save by buying the bag with the lower unit price?

Answer by stanbon(57361)

A 50-lb bag of Glossy Coat Horse Feed costs $23.50. A 25-lb bag costs $15.50. How much money per pound would you save by buying the bag with the lower unit price?
Glossy Coat Unit Price
50 lb bag: 23.50/50 = 47 cents per lb.
---------------------------------
25 lb bag: 15.50/25 = 62 cents per lb.
================================================
Savings if you buy the 50 lb. bag: 62-47 = 15 cents per lb.
===================================================================
Cheers,
Stan H.